OpenStudy (anonymous):

Quick question: sqrt100y + sqrt25y - sqrt64y sqrt100= 10 sqrt25= 5 sqrt 64 = 8 sqrt(y) +sqrt(y) -sqrt(y) When it comes to adding square roots of variables.. what's the rule? I forgot & I can't go back in the lesson without losing my current answers.

OpenStudy (aravindg):

When it comes to adding square roots of variables?? here you have perfect squares so just add the roots

OpenStudy (agent0smith):

\[10\sqrt y + 5 \sqrt y -8 \sqrt y = \] add them as you would any variables, pull a common factor... \[10\sqrt y + 5 \sqrt y -8 \sqrt y = (10 +5 -8) \sqrt y\]